# AnomalyGFM: Graph Foundation Model for Zero/Few-shot Anomaly Detection

### v1

@@ Abstract.
Graph anomaly detection (GAD) aims to identify abnormal nodes that differ from the majority of the nodes in a graph, which has been attracting significant attention in recent years. Existing generalist graph models have achieved remarkable success in different graph tasks but struggle to generalize to the GAD task. This limitation arises from their difficulty in learning generalized knowledge for capturing the inherently infrequent, irregular and heterogeneous abnormality patterns in graphs from different domains. To address this challenge, we propose AnomalyGFM, a GAD-oriented graph foundation model that supports zero-shot inference and few-shot prompt tuning for GAD in diverse graph datasets. One key insight is that graph-agnostic representations for normal and abnormal classes are required to support effective zero/few-shot GAD across different graphs. Motivated by this, AnomalyGFM is pre-trained to align data-independent, learnable normal and abnormal class prototypes with node representation residuals (i.e., representation deviation of a node from its neighbors). The residual features essentially project the node information into a unified feature space where we can effectively measure the abnormality of nodes from different graphs in a consistent way. This provides a driving force for the learning of graph-agnostic, discriminative prototypes for the normal and abnormal classes, which can be used to enable zero-shot GAD on new graphs, including very large-scale graphs. If there are few-shot labeled normal nodes available in the new graphs, AnomalyGFM can further support prompt tuning to leverage these nodes for better adaptation. Comprehensive experiments on 11 widely-used GAD datasets with real anomalies, covering social networks, finance networks, and co-review networks, demonstrate that AnomalyGFM significantly outperforms state-of-the-art competing methods under both zero- and few-shot GAD settings. Our code will be available at https://github.com/mala-lab/AnomalyGFM.
Anomaly Detection, Graph Neural Network, Graph Anomaly Detection, Foundation Model
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: the 31th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 3–7, 2025; Toronto, Canadaprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Graph anomaly detection (GAD) aims to identify anomalous nodes within a graph that deviate significantly from others. It has found broad applications in areas such as financial fraud detection, transaction analysis, and social networks (Akoglu et al., 2015; Ma et al., 2021; Pang et al., 2021; Qiao et al., 2024). Most existing GAD methods, either supervised or unsupervised methods, follow a paradigm of training and inference on the same graph, making the trained models struggle to generalize to new/unseen graphs due to the distribution shift between the training and testing graphs (Qiao and Pang, 2023a; Liu et al., 2021b; Ding et al., 2019; Huang et al., 2022). Further, this type of methods can become inapplicable in applications where graph data is not accessible during training due to data privacy or other data access issues. Generalist/foundation models for graphs have achieved remarkable progress in tackling these challenges in general tasks such as node classification, graph classification, and link prediction (Sun et al., 2023; Li et al., 2024; Zhao et al., 2024; Zi et al., 2024; Fang et al., 2024), but they are difficult to generalize the GAD task. The main reason is that the generalized knowledge learned in these generalist models through generic pre-training or tuning methods incline to the prevalent regular patterns, which differ significantly from the inherently infrequent, irregular and heterogeneous abnormality patterns across the graphs (Qiao et al., 2024; Liu et al., 2024; Lin et al., 2024).
Figure 1. Given a set of graph datasets consisting of normal nodes (blue) and abnormal nodes (red) as in (a), AnomalyGFM aims to learn graph-agnostic prototypes for normal and abnormal nodes in a projected residual feature space. Compared to the original feature representation space where normal and abnormal patterns are irregularly distributed and differ significantly in different datasets as in (b), the residual feature space learned by AnomalyGFM in (c) is universally effective in separating anomalies from normal nodes in different graphs. In (b) and (c), AnomalyGFM is trained on Facebook (Xu et al., 2022) and evaluated directly on Amazon (Dou et al., 2020) and Question (Platonov et al., 2023); both feature spaces are learned by AnomalyGFM.
@@
To fill this gap, in this paper, we introduce AnomalyGFM, a GAD-oriented graph foundation model (GFM) that is capable of addressing both zero-shot and few-shot scenarios effectively for GAD. One key insight is that to avoid overfitting the pre-training graph data, graph-agnostic representations for normal and abnormal classes are required to support effective zero/few-shot GAD across different graphs. Motivated by this, AnomalyGFM is pre-trained to align data-independent, learnable normal and abnormal class prototypes with node representation residuals (i.e., representation deviation of a node from its neighboring nodes). Compared to the original feature representation space where normal and abnormal patterns are irregularly and heterogeneously distributed in different datasets (see Fig. 1 (b)), the residual features essentially project the node information into a unified feature space where we can effectively measure the abnormality of nodes from different graphs in a consistently identical way, as shown in Fig. 1 (c). This is because normal nodes, regardless of which graphs they are from, are expected to have small residual/deviation from its neighboring nodes, whereas such residuals for abnormal nodes are large (Qiao and Pang, 2023a; Liu et al., 2024). By learning to align such residual features with the learnable normal and abnormal class prototypes, AnomalyGFM distills the residual feature-based anomaly discriminability into the two graph-agnostic prototypes, facilitating strong generalization to GAD in new graphs without any further tuning/training. Furthermore, if there are few-shot labeled normal nodes available in the new graphs, AnomalyGFM can further support graph prompt tuning to leverage these nodes for better adaptation.
@@
  • A comprehensive benchmark on both zero-shot and few-shot settings using 11 real-world GAD datasets is established, on which i) AnomalyGFM performs significantly better the state-of-the-art unsupervised, supervised, and generalist GAD methods, and ii) AnomalyGFM can scale up to very large graphs.
@@ 2.1. Graph Anomaly Detection
Existing GAD methods can be generally classified into unsupervised and supervised methods (Ma et al., 2021; Qiao et al., 2024), depending on the assumption regarding the presence of normal and anomaly labels. The fully supervised approaches have recently achieved remarkable progress by leveraging both labeled normal and abnormal nodes (Tang et al., 2022; Liu et al., 2021a; Gao et al., 2023; Wang et al., 2023). These methods are primarily designed to improve message aggregation in graph neural networks (GNNs) by reducing heterophilic edges from both spectral and spatial perspectives, effectively mitigating the over-smoothing problem in GAD (Tang et al., 2023; Qiao et al., 2024). Fully supervised methods are highly effective in GAD as they treat the task as an imbalanced classification problem. However, the requirement for both normal and abnormal labels significantly hinders their applications to scenarios where labeled nodes are difficult to obtain.
Unsupervised methods, which typically assume the absence of both normal and abnormal labels, have garnered significant attention due to their more practical setting assumption on data labels. They generally incorporate some conventional techniques such as reconstruction (Ding et al., 2019), one-class classification (Wang et al., 2021; Zhou et al., 2021; Qiao and Pang, 2023a), contrastive learning (Liu et al., 2021b; Pan et al., 2023), and adversarial learning (Ding et al., 2021) into graph learning to capture the normal patterns within the graph and then assign an anomaly score for each node based on its deviation from the normal patterns. However, these methods still follow the paradigm of training and inference on the same graph, making them struggle to generalize to new/unseen graphs due to the distribution shift between training and testing set.
@@ 3.1. Preliminaries
Few-shot GAD. In the few-shot GAD setting, the model is fine-tuned on $\mathcal{G}_{Train}$ with small labeled nodes and then applied to its test graph $\mathcal{G}_{Test}$. We denote the labeled node in the test graph as $\mathcal{V}_{l}^{Test}$ while the unlabeled nodes in the test graph are denoted as $\mathcal{V}_{u}^{Test}=\mathcal{V}^{Test}\setminus\mathcal{V}_{l}^{Test}$. Here we only consider the case of having few-shot labeled normal nodes since the normal node labels are easier/less costly to obtain than the abnormal ones (Liu et al., 2024).
@@
where ${\bf{H}}^{(\ell)}\in\mathbb{R}^{N\times h^{(l)}}$, ${\bf{H}}^{(\ell-1)}\in\mathbb{R}^{N\times h^{(l-1)}}$ are the representations of all $N$ nodes in the ${(\ell)}$-th layer and ${(\ell-1)}$-th layer, respectively, $h^{(l)}$ is the dimension of ${(\ell)}$-th layer, ${\bf{W}}^{(\ell)}$ are the learnable parameters, and ${\bf{H}}^{(0)}=\tilde{\mathbf{X}}$. ${{{\bf{H}}}=\left\{{{{\bf{h}}_{1}},{{\bf{h}}_{2}},\ldots,{{\bf{h}}_{N}}}\right\}}$ is a set of representations of $N$ nodes in the last GNN layer.
@@ 3.2. Approach Overview
AnomalyGFM aims to distill discriminative node residual features in a unified feature space into two graph-agnostic prototypes for the normal and abnormal classes, facilitating the establishment of a foundation model that supports effective zero/few-shot GAD across graphs. Specifically, AnomalyGFM consists of pre-training based on the prototype alignment, few-shot graph prompt fine-tuning, and zero-shot/few-shot inference, as shown in Fig. 2. In the pre-training step, we aim to align the data-independent and learnable normal class and abnormal class prototypes with the nodes’ residua features to achieve the distillation of graph-agnostic discriminative class-level representations. In the few-shot graph prompt fine-tuning, the labeled normal nodes’ residual features are utilized to guide a better adaptation of the normal class prototype. In the inference step, we devise zero-shot inference on new graphs, including subgraph-based inference for very large-scale graphs, and few-shot inference, both of which compute anomaly scores by measuring the similarity between the residual features of nodes and the two class prototypes.
Figure 2. Overview of AnomalyGFM. (a) During the pre-training, AnomalyGFM aims to align two learnable class-level prototypes with the representation residuals of normal and abnormal nodes to learn discriminative graph-agnostic prototypes on an auxiliary labeled graph. (b) During inference, the anomaly score is determined in a consistent way using similarity between node representation residual and prototypes where the all learnable parameters are frozen. Given a target graph with few-shot labeled normal nodes, these nodes are used to better adapt the pre-trained normal class prototypes to the target graph through a one-class prompt tuning method.
@@ 3.3. Pre-training via Prototype Alignment in a Unified Feature Space
Graph-agnostic Prototype Alignment. To exploit the anomaly discriminability in node representation residuals, we learn two data-independent learnable prototypes for the two classes that align with the normal and abnormal residual features, respectively. The alignment essentially distills the residual feature-based anomaly discriminability into the two graph-agnostic prototypes, facilitating a strong generalization to GAD in new graphs. Formally, we integrate two learnable mapping functions ${\bf p}_{n}=\Phi({\bf{z}}_{n};\Theta_{n})$, where ${\bf{z}}_{n},{\bf p}_{n}\in\mathbb{R}^{T}$ and ${\bf p}_{a}=\Phi({\bf{z}}_{a};\Theta_{a})$, where ${\bf{z}}_{a},{\bf p}_{a}\in\mathbb{R}^{T}$ to map their initializations (e.g., prototypes initialized with vectors drawn from a prior distribution such as Gaussian distribution) into the prototype space to serve as normal class prototype $\mathbf{p}_{n}$ and abnormal class prototype $\mathbf{p}_{a}$, respectively. Both $\Theta_{n}$ and $\Theta_{a}$ are the learnable parameters, and $T$ is the dimension of the learnable prototype which is set to the dimension of node representation. The Gaussian prior distributions ${\bf{z}}_{n}$ and ${\bf{z}}_{a}$ are uniformly defined as $\mathcal{N}(\mu,\sigma)$ for the normal and abnormal classes, where ${\mu}$ is the mean of the distribution and $\sigma$ is its covariance.
@@
(4) ${{L_{Alignment{\rm{}}}}=\sum\limits_{i=1}^{\left|{{{\mathcal{V}}}}\right|}{{I_{{y_{i=0}}}}}\left\|{{{\bf{r}}_{i}}-{{\bf{p}}_{n}}}\right\|_{2}^{2}+{I_{{y_{i=1}}}}\left\|{{{\bf{r}}_{i}}-{{\bf{p}}_{a}}}\right\|_{2}^{2}}.$
This distills the graph-agnostic discriminability in the residual feature space into the two prototypes. To enhance the anomaly discriminability in the node representation residual, we utilize a standard binary classification loss $L_{BCE}$ to learn discriminative representations for the normal and abnormal nodes in the original node representation space:
@@
where ${{p_{i}}={f_{\theta}}\left({{\bf h}_{i}}\right)}$ and $f_{\theta}(\cdot)$ is a Multi-Layer Perceptron (MLP) classifier with training parameters $\theta$, which maps the node representation to the probability for a node being an anomaly. Finally, we jointly optimize the ${L_{Alignment}}$ and $L_{BCE}$ via the following total loss function:

### v2

@@ Abstract.
Graph anomaly detection (GAD) aims to identify abnormal nodes that differ from the majority of the nodes in a graph, which has been attracting significant attention in recent years. Existing generalist graph models have achieved remarkable success in different graph tasks but struggle to generalize to the GAD task. This limitation arises from their difficulty in learning generalized knowledge for capturing the inherently infrequent, irregular and heterogeneous abnormality patterns in graphs from different domains. To address this challenge, we propose AnomalyGFM, a GAD-oriented graph foundation model that supports zero-shot inference and few-shot prompt tuning for GAD in diverse graph datasets. One key insight is that graph-agnostic representations for normal and abnormal classes are required to support effective zero/few-shot GAD across different graphs. Motivated by this, AnomalyGFM is pre-trained to align data-independent, learnable normal and abnormal class prototypes with node representation residuals (i.e., representation deviation of a node from its neighbors). The residual features essentially project the node information into a unified feature space where we can effectively measure the abnormality of nodes from different graphs in a consistent way. This provides a driving force for the learning of graph-agnostic, discriminative prototypes for the normal and abnormal classes, which can be used to enable zero-shot GAD on new graphs, including very large-scale graphs. If there are few-shot labeled normal nodes available in the new graphs, AnomalyGFM can further support prompt tuning to leverage these nodes for better adaptation. Comprehensive experiments on 11 widely-used GAD datasets with real anomalies, covering social networks, finance networks, and co-review networks, demonstrate that AnomalyGFM significantly outperforms state-of-the-art competing methods under both zero- and few-shot GAD settings. Code is available at https://github.com/mala-lab/AnomalyGFM.
Anomaly Detection, Graph Neural Networks, Graph Anomaly Detection, Graph Foundation Models
journalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3736843isbn: 979-8-4007-1454-2/2025/08conference: the 31th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 3–7, 2025; Toronto, Canadaccs: Computing methodologies Machine learning
@@ 1. Introduction
Graph anomaly detection (GAD) aims to identify anomalous nodes within a graph that deviate significantly from others. It has found broad applications in areas such as financial fraud detection, transaction analysis, and social networks (Akoglu et al., 2015; Ma et al., 2021; Pang et al., 2021; Qiao et al., 2024). Most existing GAD methods, either supervised or unsupervised methods, follow a paradigm of training and inference on the same graph, making the trained models struggle to generalize to new/unseen graphs due to the distribution shift between the training and testing graphs (Qiao and Pang, 2023; Liu et al., 2021b; Ding et al., 2019). Further, this type of methods can become inapplicable in applications where graph data is not accessible during training due to data privacy or other data access issues. Generalist/foundation models for graphs have achieved remarkable progress in tackling these challenges in general tasks such as node classification, graph classification, and link prediction (Sun et al., 2023; Li et al., 2024; Zhao et al., 2024; Zi et al., 2024; Fang et al., 2024), but they are difficult to generalize to the GAD task. The main reason is that the generalized knowledge learned in these generalist models through generic pre-training or tuning methods inclines to the prevalent regular patterns, which differ significantly from the inherently infrequent, irregular, and heterogeneous abnormality patterns across the graphs (Qiao et al., 2024; Liu et al., 2024; Lin et al., 2024).
Figure 1. Given a set of graph datasets consisting of normal nodes (blue) and abnormal nodes (red) as in (a), AnomalyGFM aims to learn graph-agnostic prototypes for normal and abnormal nodes in a projected residual feature space. Compared to the original feature space where normal and abnormal patterns are irregularly distributed and differ significantly in different datasets as in (b), the residual feature space learned by AnomalyGFM in (c) is universally effective in separating anomalies from normal nodes in different graphs. In (b) and (c), AnomalyGFM is trained on Facebook (Xu et al., 2022) and evaluated directly on Amazon (Dou et al., 2020) and Question (Platonov et al., 2023); both feature spaces are learned by AnomalyGFM.
@@
To fill this gap, in this paper, we introduce AnomalyGFM, a GAD-oriented graph foundation model (GFM) that is capable of addressing both zero-shot and few-shot scenarios effectively for GAD. One key insight is that to avoid overfitting the pre-training graph data, graph-agnostic representations for normal and abnormal classes are required to support effective zero/few-shot GAD across different graphs. Motivated by this, AnomalyGFM is pre-trained to align data-independent, learnable normal and abnormal class prototypes with node representation residuals (i.e., representation deviation of a node from its neighboring nodes). Compared to the original feature space where normal and abnormal patterns are irregularly and heterogeneously distributed in different datasets (see Fig. 1 (b)), the residual features essentially project the node information into a unified feature space where we can effectively measure the abnormality of nodes from different graphs in a consistently identical way, as shown in Fig. 1 (c). This is because normal nodes, regardless of which graphs they are from, are expected to have small residual/deviation from its neighboring nodes, whereas such residuals for abnormal nodes are large (Qiao and Pang, 2023; Liu et al., 2024). By learning to align such residual features with the learnable normal and abnormal class prototypes, AnomalyGFM distills the residual feature-based anomaly discriminability into the two graph-agnostic prototypes, facilitating strong generalization to GAD in new graphs without any further tuning/training. Furthermore, if there are few-shot labeled normal nodes available in the new graphs, AnomalyGFM can further support graph prompt tuning to leverage these nodes for better adaptation.
@@
  • A comprehensive benchmark on both zero-shot and few-shot settings using 11 real-world GAD datasets is established, on which i) AnomalyGFM performs significantly better than the state-of-the-art unsupervised, supervised, and generalist GAD methods, and ii) AnomalyGFM can scale up to very large graphs.
@@ 2.1. Graph Anomaly Detection
Existing GAD methods can be generally classified into unsupervised and supervised methods (Ma et al., 2021; Qiao et al., 2024), depending on the assumption regarding the presence of normal and anomaly labels. The fully supervised approaches have recently achieved remarkable progress by leveraging both labeled normal and abnormal nodes (Tang et al., 2022; Liu et al., 2021a; Gao et al., 2023; Wang et al., 2023). These methods are primarily designed to improve message aggregation in graph neural networks (GNNs) by reducing heterophilic edges from both spectral and spatial perspectives, effectively mitigating the over-smoothing problem in GAD (Tang et al., 2023; Qiao et al., 2024). Fully supervised methods are effective in GAD as they treat the task as an imbalanced classification problem. However, the requirement for both normal and abnormal labels significantly hinders their applications to scenarios where labeled nodes are difficult to obtain.
Unsupervised methods, which typically assume the absence of both normal and abnormal labels, have garnered significant attention due to their more practical setting assumption on data labels. They generally incorporate some conventional techniques such as reconstruction (Ding et al., 2019), one-class classification (Wang et al., 2021; Zhou et al., 2021; Qiao and Pang, 2023), contrastive learning (Liu et al., 2021b; Pan et al., 2023), and adversarial learning (Ding et al., 2021) into graph learning to capture the normal patterns within the graph and then assign an anomaly score for each node based on its deviation from the normal patterns. However, these methods still follow the paradigm of training and inference on the same graph, making them struggle to generalize to new/unseen graphs due to the distribution shift between training and testing set.
@@ 3.1. Preliminaries
Few-shot GAD. In the few-shot GAD setting, the model is fine-tuned on $\mathcal{G}_{Test}$ with small labeled nodes and then applied to the rest of the unlabelled nodes in the test graph $\mathcal{G}_{Test}$. We denote the labeled node in the test graph as $\mathcal{V}_{l}^{Test}$ while the unlabeled nodes in the test graph are denoted as $\mathcal{V}_{u}^{Test}=\mathcal{V}^{Test}\setminus\mathcal{V}_{l}^{Test}$.
@@
where ${\bf{H}}^{(\ell)}\in\mathbb{R}^{N\times h^{(l)}}$, ${\bf{H}}^{(\ell-1)}\in\mathbb{R}^{N\times h^{(l-1)}}$ are the representations of all $N$ nodes in the ${(\ell)}$-th layer and ${(\ell-1)}$-th layer, respectively, $h^{(l)}$ is the dimension of ${(\ell)}$-th layer, ${\bf{W}}^{(\ell)}$ are the learnable parameters, and ${\bf{H}}^{(0)}=\tilde{\mathbf{X}}$. ${{{\bf{H}}}=\left\{{{{\bf{h}}_{1}},{{\bf{h}}_{2}},\ldots,{{\bf{h}}_{N}}}\right\}}$ is a set of representations of $N$ nodes in the last GNN layer. In this paper, we adopt a 2-layer GCN to model the graph due to its efficient and simple architecture.
@@ 3.2. Approach Overview
AnomalyGFM aims to distill discriminative node residual features in a unified feature space into two graph-agnostic prototypes for the normal and abnormal classes, facilitating the establishment of a foundation model that supports effective zero/few-shot GAD across graphs. Specifically, AnomalyGFM consists of pre-training based on the prototype alignment, few-shot graph prompt fine-tuning, and zero-shot/few-shot inference, as shown in Fig. 2. In the pre-training step, we aim to align the data-independent and learnable normal class and abnormal class prototypes with the nodes’ residual features to achieve the distillation of graph-agnostic discriminative class-level representations. In the few-shot graph prompt fine-tuning, the labeled normal nodes’ residual features are utilized to guide a better adaptation of the normal class prototype. In the inference, we devise zero-shot inference on new graphs, including subgraph-based inference for very large-scale graphs, and few-shot inference, both of which compute anomaly scores by measuring the similarity between the residual features of nodes and the two class prototypes.
Figure 2. Overview of AnomalyGFM. (a) During the pre-training, AnomalyGFM aims to align two learnable class-level prototypes with the representation residuals of normal and abnormal nodes to learn discriminative graph-agnostic prototypes on an auxiliary labeled graph. (b) During inference, the anomaly score is determined in a consistent way using similarity between node representation residual and prototypes where all learnable parameters are frozen. Given a target graph with few-shot labeled normal nodes, these nodes are used to better adapt the pre-trained normal class prototypes to the target graph through a one-class prompt tuning method.
@@ 3.3. Pre-training via Prototype Alignment in a Unified Feature Space
Graph-agnostic Prototype Alignment. To exploit the anomaly discriminability in node representation residuals, we learn two data-independent learnable prototypes for the two classes that align with the normal and abnormal residual features, respectively. The alignment essentially distills the residual feature-based anomaly discriminability into the two graph-agnostic prototypes, facilitating a strong generalization to GAD in new graphs. Formally, we integrate two learnable mapping functions ${\bf p}_{n}=\Phi({\bf{z}}_{n};\Theta_{n})$, where ${\bf{z}}_{n},{\bf p}_{n}\in\mathbb{R}^{T}$ and ${\bf p}_{a}=\Phi({\bf{z}}_{a};\Theta_{a})$, where ${\bf{z}}_{a},{\bf p}_{a}\in\mathbb{R}^{T}$ to map their initializations (e.g., prototypes initialized with vectors drawn from a prior distribution such as Gaussian distribution) into the prototype space to serve as normal class prototype $\mathbf{p}_{n}$ and abnormal class prototype $\mathbf{p}_{a}$, respectively. Both $\Theta_{n}$ and $\Theta_{a}$ are the learnable parameters, and $T$ is the dimension of the learnable prototype, which is set to match the dimensionality of the node representation $d$. The Gaussian prior distributions ${\bf{z}}_{n}$ and ${\bf{z}}_{a}$ are uniformly defined as $\mathcal{N}(\mu,\sigma)$ for the normal and abnormal classes, where ${\mu}$ is the mean of the distribution and $\sigma$ is its covariance.
@@
(4) ${{L_{Alignment{\rm{}}}}=\sum\limits_{i=1}^{\left|{{{\mathcal{V}}}}\right|}{{I_{{y_{i=0}}}}}\left\|{{{\bf{r}}_{i}}-{{\bf{p}}_{n}}}\right\|_{2}^{2}+{I_{{y_{i=1}}}}\left\|{{{\bf{r}}_{i}}-{{\bf{p}}_{a}}}\right\|_{2}^{2}}\,,$
where ${I_{{y_{i=0}}}}$ is an indicator function, which takes the value of 1 when the condition is satisfied. This distills the graph-agnostic discriminability in the residual feature space into the two prototypes. To enhance the anomaly discriminability in the node representation residual, we utilize a standard binary classification loss $L_{BCE}$ to learn discriminative representations for the normal and abnormal nodes in the original node representation space:
@@
where ${{p_{i}}={f_{\theta}}\left({{\bf h}_{i}}\right)}$ and $f_{\theta}(\cdot)$ is a Multi-Layer Perceptron (MLP) classifier with training parameters $\theta$, which maps the representation to the probability for a node being an anomaly. Finally, we jointly optimize the ${L_{Alignment}}$ and $L_{BCE}$ via the following total loss :